The formule D=0.054x^2+0.058x describes the distance in feet D that it takes to stop a vehicle traveling x miles per hour on dry pavement.
a. How fast can you drive if you wish to be able to stop your car within 65 feet?
b.On black ice, a truck's stopping distance is 3 times its stopping distance on dry pavement. A truck traveling 20 miles per hour applies the brakes, on black ice, at a distance of 65 feet in front of a rubber traffic cone. Will the truck hit the cone?

Answer by stanbon(75887)

The formule D=0.054x^2+0.058x describes the distance in feet D that it takes to stop a vehicle traveling x miles per hour on dry pavement.
------
a. How fast can you drive if you wish to be able to stop your car within 65 feet?
Solve: 0.054x^2+0.058x = 65
---
Graph to find "x":
Ans: 34.16 mph
------------------------------------
b.On black ice, a truck's stopping distance is 3 times its stopping distance on dry pavement. A truck traveling 20 miles per hour applies the brakes, on black ice, at a distance of 65 feet in front of a rubber traffic cone. Will the truck hit the cone?
---
Normal stopping distance:
D(20) = 0.54*20^2+0.058*20 = 22.76 ft.
-----
On black ice: 3*22.76 = 68.28 ft
====================
Cheers,
Stan H.
====================